OpenStudy (anonymous):

Helllllpppp!!!!!!!!!!!1 A 120-foot-long cable of diameter 5 inches is submerged in seawater. Because of corrosion, the surface area of the cable decreases at the rate of 760 in2 per year. Express the diameter d of the cable as a function of time t (in years). (Disregard corrosion at the ends of the cable.)

OpenStudy (anonymous):

d(t) = 600*pi - 760t? verify

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

im thinking 5(600pi - 760t/500pi)

OpenStudy (anonymous):

make the last one 600 pi

OpenStudy (thomas5267):

I calculated \[\frac{152 \text{in}^2}{\pi(\text{year})}\] but I am not sure as I am new to calculus.

OpenStudy (anonymous):

is this Calculus or PreCalculus?

OpenStudy (anonymous):

it's pre

OpenStudy (anonymous):

i think we need to use a surface of right cylinder

OpenStudy (anonymous):

yep

OpenStudy (anonymous):